OpenStudy (anonymous):

Evaluate the indefinite integral of: Sin 2x / 1 + cos^(2)(x) * dx

hero (hero):

If you let: \(u = cos(x)\\du = -sin(u) \\ dx = \frac{-du}{sin(x)}\) Then the integral becomes: \[\int -\frac{1}{u^2 + 1} du\]

hero (hero):

Which \[\frac{1}{u^2 + 1} = \tan^{-1}(u) = \tan^{-1}(\cos(x))\]